

********************************************************************
UARS WEEKLY STATUS REPORT - 08 MAR 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was turned off on March 7 for the yaw around to save power.
Otherwise, the instrument performed its daily mid-resolution and
MgII solar scans during periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothig received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime winds
when power is available.
The HRDI Operations Team successfully converted our STOL procedures to CECIL.
We extend a special thanks to Greg Haun from GSFC for running his
autoconverter and for much help on a few undocumented lines that had to be
converted manually. We would also like to thank Lynn Herring from SUSIM for
providing some helpful insight into the process.
The team has also succeeded recently in updating our data archiving
system. In recent years, we had stored data on 10 or so disks spread
over 5 machines in our OpenVMS Alpha cluster. To facilitate migration
off OpenVMS and for much increased reliability, we moved our archive to
a Raidtec SNAZ6 network-attached storage system. We configured it to
have 288Gb data space, RAID 5 redundancy, and a hot spare drive. (This
means we could lose 2 drives without losing any data!)
We can still access our data through OpenVMS because we have
mounted the new server through NFS and all of our existing systems work
without difficulty. Plus, we can access our data directly through MS
Windows machines since the server runs Microsoft networking as well. The new
system has been well worth the time and capital spent.
This whole process was expedited due to our recent loss of one of our
archive disks, and our inability to recover all our files from backup tapes.
Thanks to Frank Huang and Mike Anderson and anyone else who helped us get our
files from the old CDHF2 archive system.
***********
WINDII
***********
WINDII is preparing to operate March 13 to April 8, in coordination
with the TIDI instrument on TIMED.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 February 2002 (Orbit 57153, GMT Day 0547)
through 01 March 2002 (Orbit 57258, GMT Day 060).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-None this report period.
The beta angle ranged between the angles of 44.31 degrees to 28.67
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 08 March 2002 (DOY 067).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 24.32 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.94 to 1.31 degrees C. Additional
battery performance data for each battery on 23 February 2002 and 01
March 2002 are provided in Appendix D.
The clock error ranged from -10.8 to 11.7 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 8
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 0
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/002 - 2287. 498845
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Eclipse testing remained the focal point this past week on UARS. The
integration of the new system has reached its final month and only a
few weeks are left before the April 1st deadline. The FOT continues
to test all of the Eclipse functionality against the UTTS and against
the spacecraft. Some errors have been found but the Raytheon software
development team has done an excellent job of turning around the fixes
for these problems. Thus the integration still remains on schedule.
A test from Eclipse to the spacecraft occurred on 27 February 2002.
The test performed was a partial table load without checksum of table
15. The test was successful and the table load was verified via
telemetry as well as via a dump and compare on the Eclipse system.
More tests including ATC and Ephemeris loads will take place next
week.
The FOT is also trying very hard to finalize the conversion of all
STOL procedures to the new CECIL format. All the procedures are
scheduled to be sent out to the instrument teams for their review and
approval. The FOT would like to thank them in advance for their quick
response.
MEETINGS
The FOT attended a UARS Eclipse Working Meeting on Thursday, February
28, 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 08 March 2002 (DOY 067). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 01 March 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
57164 054/1748 02-062 OBC Clk Adj 0.00 to 20.6 msec/day
57186 056/0504 02-063 OBC Clk Adj 20.6 to 0.00 msec/day
57225 058/1844 02-064 OBC Memory Monitor (ECLIPSE)
57225 058/1923 02-065 OBC Dump Table 15
57164 054/1748 02-062 OBC Clk Adj 0.00 to 20.6 msec/day
57186 056/0504 02-063 OBC Clk Adj 20.6 to 0.00 msec/day
57225 058/1844 02-064 OBC Memory Monitor (ECLIPSE)
57225 058/1923 02-065 OBC Dump Table 15
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 01 March 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
_____________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 01 March 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
57164 054/1748 Clk Adj 0.00 to 20.6 msec/day
57186 056/0504 Clk Adj 20.6 to 0.00 msec/day
57164 054/1748 Clk Adj 0.00 to 20.6 msec/day
57186 056/0504 Clk Adj 20.6 to 0.00 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 01 March 2002
23 February 2002 - GMT Day 054 Beta = 44.31 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -50.4/-67.2 -3.16/-4.38 +0.0/ -0.0 10.72 0.000 0.00
2 +11.2/-44.8 +5.14/+3.82 +25.6/-14.4 26.08 1.040 19.92
3 +22.4/-22.4 +3.82/+2.85 +25.2/-14.4 26.08 1.030 19.69
01 March 2002 - GMT Day 060 Beta = 28.67 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 -44.8/-61.6 -3.16/-4.38 +0.0/ -0.0 10.72 0.000 0.00
2 +11.2/-84.0 +6.14/+5.14 +29.6/-15.2 24.48 1.020 21.90
3 +22.4/-50.4 +5.14/+4.15 +28.8/-15.2 24.32 1.020 21.67
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 01 March 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-005 FOT see list Feb 02 14170 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode